Use the First Derivative Test to find the relative extrema of the function, if they exist. f(x) = x^4 - 2x^2 + 5
Answer #1

fun 1 Pg - 1 Solution first Derivative test for nelestive extremas suppore that cis critical point of con timour of f, cnd fi60, 110-27 EMO) -4 + 2 1) 12 ए (-0, -1) F(-2) = 46-233-46-2) = -32 +9 = -24 <0 (-1,0) f(-1) 46-312) 3 - 414): :(0,1) 2 / 0 f